x-intercept 2, y- intercept 1/3 [write an equation for this line]
Answer by TimothyLamb(4379)

Use the intercept form: x/a + y/b = 1
---
where a is the x-intercept and b is the y-intercept:
a = 2
b = (1/3)
---
x/a + y/b = 1
x/2 + y/(1/3) = 1
x/2 + 3y = 1
3y = -x/2 + 1
---
Answer:
y = (-1/6)x + (1/3)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php